# Investigating the role of mutual information in the Page curve for a functional renormalization group improved Schwarzschild black hole

### v1

@@ IV Conclusion
We will now summerize our findings. In this paper, we have investigated the crucial role played by the mutual information (in the after Page time scenario) between the subsystems (relevant regions where the matter fields are localized) $B_{+}$ and $B_{-}$ in obtaining the correct time-evolution of fine-grained entropy of Hawking radiation, as suggested by Page. This was done by utilizing a proposal first given in Saha et al. (2022) and later followed in Roy Chowdhury et al. (2022, 2023). This we have done for a quantum corrected or more precisely, a functional renormalization group improved two-sided eternal Schwarzschild black hole Bonanno and Reuter (2000). The motivation for considering such non-trivial geometry lies in the fact that this solution fimrly represents the essence of an “effective” quantum theory of gravity as it arises from asymptotically safe effective average action $\Gamma_{k}[g_{\mu\nu}]$ which can correctly describe all gravitational phenomena, including the effects of all loops Reuter (1998). Further, the mentioned action leads to a running nature for the usual Newton’s gravitational constant which discriminates this geometry from the standard Schwarzschild black hole solution. On the other hand, the mentioned proposal regarding mutual information can be justified form the following facts. One can note that just after the Page time $t_{P}$, replica wormhole saddle point dominates and the island starts to contribute significantly and one has to make use of the formula given in eq.(1), in order to compute the fine-grained entropy of Hawking radiation correctly where the first term provides us a geometric contribution and the second term quantifies the matter field contribution. Further, based upon the pure state property one can write down $S_{vN}(I\cup R)=S_{vN}(I\cup R_{+}\cup R_{-})=S_{vN}(B_{+}\cup B_{-})$. In earlier works in this direction it was observed that by using the late time approximated result of $S_{vN}(B_{+}\cup B_{-})$, that is, $S_{vN}(B_{+}\cup B_{-})\sim S_{vN}(B_{+})+S_{vN}(B_{-})$ along with the extremized values of the island parameters ($t_{a}\sim t_{b}$ and $a\sim r_{H}$), one gets the desired time-independent result for $S(R)$ corresponding to a two-sided eternal black hole geometry in the after Page time scenario. However, the approximation result of $S_{vN}(B_{+}\cup B_{-})\sim S_{vN}(B_{+}\cup B_{-})$ deals with the fact that one has to ignore time-dependent terms $\sim e^{\frac{2\pi t_{b,a}}{\beta^{q}}}$ and if not neglected, obtaining a time-independent expression for $S(R)$ would not be possible. This in turn means one has to ignore certain time-dependent terms in order to have a desired time-independent result. We resolved this ambiguity in Saha et al. (2022); Roy Chowdhury et al. (2022, 2023) by demanding that the mutual information between $B_{+}$ and $B_{-}$ vanishes exactly, purely due to the emergence of island just after the Page time. We follow the same path here. This justified demand of vanishing mutual correlation leads to the associated condition $t_{a}-t_{b}=|r^{*}(a)-r^{*}(b)|$ and by utilizing this condition we have finally obatined the time-independent expression of $S_{vN}(I\cup R)$. This is a very important result as it helps one to overcome the mentioned ambiguity of neglecting certain time-dependent terms and also one obtains the needed value of $t_{a}$ without performing a extremization. Now extremizing the result of $S(R)$ with respect to the island parameter “$a$”, we have found out the position of the quantum extremal surfaces. Based upon this observations, we also propose a modified form of the island formula (given in (II)) which one may use to evaluate the fine-grained entropy of Hawking radiation for an eternal black hole. We make use of this extremized value of ``$a"$ in two different places, firstly, by using it we get the final expression for $S(R)$ and on the other hand it simplifies the condition for vanishing of mutual information in the following way $t_{a}-t_{b}\sim t_{Scr}^{q}$. This result reflects the fact that as long as $t_{a}-t_{b}<t_{Scr}^{q}$, the mutual information is non-zero and the entanglement wedge of $B_{+}\cup B_{-}$ is in its connected phase and for $t_{a}-t_{b}=t_{Scr}^{q}$, the mutual information vanishes, depicting the disconnected phase of entnaglement wedge associated to $B_{+}\cup B_{-}$. Another interseting fact is that, the scrambling time for the quantum corrected Schwarzschild solution is greater than the same for standard Schwarzschild geometry. This in turn means that the information recovery time for the quantum corrected black hole is larger than the usual Schwarzschild black hole. We have also obtained the expression of the Page time for the mentioned geometry. We have found that the Page time for the quantum corrected Schwarzschild black hole is smaller compares to that of Schwarzschild black hole. Based upon this, one may say that island formation starts earlier for the QCSBH in comparison to the Schwarzschild black hole. Finally, we also study the time-dependent behaviour of relevant mutual information, that is, $I(R_{+}:R_{-})$ in both the early and late time domain. Further, this leads us to the observation that during the early time $I(R_{+}:R_{-})$ decreases with respect to the observer’s time and at late time it increases with time. This gives us a hint that, there must exists a value of the observer’s time at which the mutual information between $R_{+}$ and $R_{-}$ vanishes and the entanglement wedge associated to $R_{+}\cup R_{-}$ gets in its disconnected phase. We compute the explicit expression for the mentioned time-scale and observe that $S(R)|_{t_{b}=t_{R}}\sim\log\beta$. This observation leads us to identify the time $t_{R}$ as the Hartman-Maldacena time.
@@ V Acknowledgements
AS would like to thank S.N. Bose National Centre for Basic Sciences for the financial support through its Advanced Postdoctoral Research Programme. ARC would like to thank the organizers of Strings 2025 held at NYU Abu Dhabi, as some of the results related to this work were presented there. The authors would also like to thank the organizers of XXVI DAE-BRNS High Energy Physics Symposium 2024 held at Banaras Hindu University, Varanasi, as some of the results related to this work were presented there.

### v2

@@ IV Conclusion
We will now summerize our findings. In this paper, we have investigated the crucial role played by the mutual information (in the after Page time scenario) between the subsystems (relevant regions where the matter fields are localized) $B_{+}$ and $B_{-}$ in obtaining the correct time-evolution of fine-grained entropy of Hawking radiation, as suggested by Page. This was done by utilizing a proposal first given in Saha et al. (2022) and later followed in Roy Chowdhury et al. (2022, 2023). This we have done for a quantum corrected or more precisely, a functional renormalization group improved two-sided eternal Schwarzschild black hole Bonanno and Reuter (2000). The motivation for considering such non-trivial geometry lies in the fact that this solution fimrly represents the essence of an “effective” quantum theory of gravity as it arises from asymptotically safe effective average action $\Gamma_{k}[g_{\mu\nu}]$ which can correctly describe all gravitational phenomena, including the effects of all loops Reuter (1998). Further, the mentioned action leads to a running nature for the usual Newton’s gravitational constant which discriminates this geometry from the standard Schwarzschild black hole solution. On the other hand, the mentioned proposal regarding mutual information can be justified form the following facts. One can note that just after the Page time $t_{P}$, replica wormhole saddle point dominates and the island starts to contribute significantly and one has to make use of the formula given in eq.(1), in order to compute the fine-grained entropy of Hawking radiation correctly where the first term provides us a geometric contribution and the second term quantifies the matter field contribution. Further, based upon the pure state property one can write down $S_{vN}(I\cup R)=S_{vN}(I\cup R_{+}\cup R_{-})=S_{vN}(B_{+}\cup B_{-})$. In earlier works in this direction it was observed that by using the late time approximated result of $S_{vN}(B_{+}\cup B_{-})$, that is, $S_{vN}(B_{+}\cup B_{-})\sim S_{vN}(B_{+})+S_{vN}(B_{-})$ along with the extremized values of the island parameters ($t_{a}\sim t_{b}$ and $a\sim r_{H}$), one gets the desired time-independent result for $S(R)$ corresponding to a two-sided eternal black hole geometry in the after Page time scenario. However, the approximation result of $S_{vN}(B_{+}\cup B_{-})\sim S_{vN}(B_{+})+S_{vN}(B_{+})$ deals with the fact that one has to ignore time-dependent terms $\sim e^{-\frac{2\pi t_{b,a}}{\beta^{q}}}$ and if not neglected, obtaining a time-independent expression for $S(R)$ would not be possible. This in turn means one has to ignore certain time-dependent terms in order to have a desired time-independent result. We resolved this ambiguity in Saha et al. (2022); Roy Chowdhury et al. (2022, 2023) by demanding that the mutual information between $B_{+}$ and $B_{-}$ vanishes exactly, purely due to the emergence of island just after the Page time. We follow the same path here. This justified demand of vanishing mutual correlation leads to the associated condition $t_{a}-t_{b}=|r^{*}(a)-r^{*}(b)|$ and by utilizing this condition we have finally obtained the time-independent expression of $S_{vN}(I\cup R)$. This is a very important result as it helps one to overcome the mentioned ambiguity of neglecting certain time-dependent terms and also one obtains the needed value of $t_{a}$ without performing a extremization. Now extremizing the result of $S(R)$ with respect to the island parameter “$a$”, we have found out the position of the quantum extremal surfaces. Based upon this observations, we also propose a modified form of the island formula (given in (II)) which one may use to evaluate the fine-grained entropy of Hawking radiation for an eternal black hole. We make use of this extremized value of ``$a"$ in two different places, firstly, by using it we get the final expression for $S(R)$ and on the other hand it simplifies the condition for vanishing of mutual information in the following way $t_{a}-t_{b}\sim t_{Scr}^{q}$. This result reflects the fact that as long as $t_{a}-t_{b}<t_{Scr}^{q}$, the mutual information is non-zero and the entanglement wedge of $B_{+}\cup B_{-}$ is in its connected phase and for $t_{a}-t_{b}=t_{Scr}^{q}$, the mutual information vanishes, depicting the disconnected phase of entnaglement wedge associated to $B_{+}\cup B_{-}$. Another interseting fact is that, the scrambling time for the quantum corrected Schwarzschild solution is greater than the same for standard Schwarzschild geometry. This in turn means that the information recovery time for the quantum corrected black hole is larger than the usual Schwarzschild black hole. We have also obtained the expression of the Page time for the mentioned geometry. We have found that the Page time for the quantum corrected Schwarzschild black hole is smaller compared to that of Schwarzschild black hole. Based upon this, one may say that island formation starts earlier for the QCSBH in comparison to the Schwarzschild black hole. Finally, we also study the time-dependent behaviour of relevant mutual information, that is, $I(R_{+}:R_{-})$ in both the early and late time domain. Further, this leads us to the observation that during the early time $I(R_{+}:R_{-})$ decreases with respect to the observer’s time and at late time it increases with time. This gives us a hint that, there must exists a value of the observer’s time at which the mutual information between $R_{+}$ and $R_{-}$ vanishes and the entanglement wedge associated to $R_{+}\cup R_{-}$ gets in its disconnected phase. We compute the explicit expression for the mentioned time-scale and observe that $S(R)|_{t_{b}=t_{R}}\sim\log\beta$. This observation leads us to identify the time $t_{R}$ as the Hartman-Maldacena time.
@@ V Acknowledgements
AS would like to thank S.N. Bose National Centre for Basic Sciences for the financial support through its Advanced Postdoctoral Research Programme. ARC would like to thank the organizers of Strings 2025 held at NYU Abu Dhabi, as some of the results related to this work were presented there. The authors would also like to thank the organizers of XXVI DAE-BRNS High Energy Physics Symposium 2024 held at Banaras Hindu University, Varanasi, as some of the results related to this work were presented there. The authors would like to thank the referees for very useful and crucial comments.